October 2, 1999 Evening shift summary

B. Cox (cox@fnal.gov)
Sun, 3 Oct 1999 00:10:10 -0500 (CDT)

Shift Summary for Evening Shift, Oct 2,/99

On Shift: Brad, Rick K.

Hours of beam available at all intensities: 8.5
Hours of beam available at nominal intensity: 8.5
Hours of E832 data (special runs not included): 8.4
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:

one fresh start to eliminate accept fraction problem

========================================================================
Runs ended this shift:

14794 : BEAM_799

=======================================================================

Tapes taken to FCC this shift

none

========================================================================
Accesses during shift:

None

========================================================================
Overview of shift:

15:30 Begin shift. Running is smooth but accept fractions are
not uniform vs. cpu. Same pehnomenon as last night during run
14789.

15:50 BobT calls. Wants to be called if there is an hour down
time.

16:00 End run and make do a fresh start to try to eliminate the
accept fraction problem. Run 14795 started.

16:02 A major new addition to the counting room in the form of
a couch appears. All that is missing is the
remote [kfrend] control and we're in business!!

16:07 Spill 3 is probably garbage due to missing ET quadrant.

16:12 Back running again

16:50 Changed the alarm limits on the RF Level 1 timing.
93.8 to 94.6 (RK)

19:00 Stan Johnson calls from Main Control. He noticed that
we had been requesting tuning to minimize the 2 Hz ripple
on the chamber current monitor. However, Greg Bock had
indicated yesterday that the 2Hz was an instrumentation problem
and could be ignored. The basis of this was that
when there was no beam (or at the end of spill), the
oscillations persisted for a while. They put Greg's
remark in their logbook and have apparently been ignoring
our requests to tune to minimize the oscillation.

We talked to Greg and informed him that smaller amplitude
appears to be correlated to higer lifetime so the monitor,
while maybe not perfect, is doing something. Greg said
he would call and correct their misimpression.

19:05 laser rms too large message, spills 133-135, recovers

20:20 Looking at the run 14794 and 14795 live time and DC current
plots from trg#live, it is seen that at approximately
15:30 the beam got considerably worse (as measured by the variation
in the DC1 current). We ask Rick Coleman to talk to Main Control
to get them to tume for better slow spill by minimizing the slow
2 Hz oscillation. Rick will look into it to see if he can
determine anything to do.

21:20 The live time has crept up to approximately 70% but without
a corresponding decrease in the fluctuations of DC1 current.
We have approximately the same live time as was recorded
in run 14794 but we have 50+-25 microamps in run 14795 compared
to 50+-10 before spill 100 in run 14794.

21:25 Rick Coleman calls back. He cannot see anything to do so we just
will have to live with it.

21:30 Intensity drifted down to less than 8E12. Call Main Control to
ask them to raise intensity. Most of the increase in live time
just due to slight decrease in beam intensity.

21:50 Spill 254 ped problem detected with laser illumination.
CsI 1662 is dead.
Try to intialize pipeline. CsI 1662 now ok

22:00 R.K. tightens L2beam/L1beam alarm window to 0.15 to 019

22:10 beam intensity back up again.

23:30 Noticed that no slow daq alarms seem to be active.
Try to generate alarms by changing limit on live time.

23:50 Finally ran 'scaler_check' from sdaq account and that
got it going again.

========================================================================
Description of Shift:

Problems with the spill structure as seen by the DC1 current
monitor (if real). Otherwise smooth running with no appreciable
downtime.

========================================================================
Follow up Notes on Previous Shift

Try to get in touch with Theo to deal with the lack of slow daq alarms.